OpenStudy (anonymous):

If a line crosses the y-axis at (0, -4) and has a slope of -2, what is the equation of the line?

OpenStudy (anonymous):

2x − y = -4 2x + y = -4 2x + y = 4 2x − y = 4

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Start by writing the equation in slope intercept form \[y = mx+b\] m = slope b = y-intercept

OpenStudy (sohailiftikhar):

use the formula for slop slop=y-y1/x-x1\[m=y-y1/x-x1\] here given point (0,-4) so x1=0 and y1=-4 and slop is -2 put these values in equation you will get -2=y+4/x -2x-y=4 or 2x+y=-4 so answer is B